# LoTLIP: Improving Language-Image Pre-training for Long Text Understanding

### v3

@@ 3 Preliminary of Language-Image Pre-training
Language-image pre-training models, i.e., CLIP [23] and LiT [33], typically consist of two parts: an image encoder and a text encoder. In the training stage, the language-image model takes image-text pairs as input, and the image encoder and text encoder are used to extract embeddings from images and texts, respectively. Then, two encoders are trained with contrastive objectives, ensuring that paired image and text embeddings are close in the embedding space, while unpaired pairs are far apart. Specifically, in the pre-training stage, a batch of images $\{I_{1},I_{2},\cdots,I_{N}\}$ and the corresponding short texts $\{T_{1},T_{2},\cdots,T_{N}\}$ are randomly sampled from the pre-training dataset. Each image and its corresponding text are treated as a positive pair, while others in the same batch are negative pairs. Then, the image encoder extracts image global feature $\mathbf{v}^{i}$ from the $i$-th image $I_{i}$ within the batch, while the text encoder obtains text feature $\mathbf{t}^{j}$ from the $j$-th text $T_{j}$. These two encoders are then optimized using contrastive loss $\mathcal{L}$, which consists of image-to-text loss $\mathcal{L}^{i2t}$ and text-to-image loss $\mathcal{L}^{t2i}$. It can be formulated as follows:
@@ 5.2 Ablation Studies
Table 2: Analyze the effectiveness of LoTLIP in language-image pre-training with long texts. The architecture of the image encoder is ViT-B/16. T2I and I2T indicate R@1 on text and image retrieval, respectively. We use 3M scale dataset for pre-training. “✓" indicates we add long texts in the training stage.
@@
T2I I2T T2I I2T T2I I2T T2I I2T Acc@1
LiT - 27.14 24.13 65.20 58.50 32.73 27.01 24.0734.20 43.76
LiT ✓ 47.96 44.92 84.97 81.70 73.66 66.73 30.06 43.52 48.87
LiT+Long-CLIP* ✓ 47.21 46.61 83.82 83.01 74.60 67.49 26.74 43.78 46.82
LoTLIP49.4647.8284.9783.3376.4969.7231.5946.5650.34
@@
T2I I2T T2I I2T T2I I2T T2I I2T Acc
0 - 47.96 44.92 84.97 81.7073.66 66.73 30.06 43.52 48.87
1 ✓ 49.57 46.55 84.97 82.6874.91 68.41 31.51 45.68 49.62
2 ✓ 49.46 47.82 84.97 83.33 76.49 69.7231.5946.56 50.34
3 ✓ 49.58 47.70 87.0984.3176.5170.20 31.60 46.48 50.36
4 ✓ 49.5848.30 86.76 84.1576.25 70.1431.88 47.7050.59
2 - 48.61 47.17 86.11 81.8676.14 69.31 31.34 47.70 49.88
2 ✓ 49.46 47.82 84.97 83.33 76.49 69.7231.5946.56 50.34
Table 4: Zero-shot evaluation of different models on long-text-image retrieval tasks. T2I and I2T indicate R@1 on text and image retrieval, respectively.

### v4

@@ 3 Preliminary of Language-Image Pre-training
Language-image pre-training models, i.e., CLIP [23] and LiT [33], typically consist of two parts: an image encoder and a text encoder. In the pre-training stage, the language-image model takes image-text pairs as input, and the image encoder and text encoder are used to extract embeddings from images and texts, respectively. Then, two encoders are trained with contrastive objectives, ensuring that paired image and text embeddings are close in the embedding space, while unpaired pairs are far apart. Specifically, a batch of images $\{I_{1},I_{2},\cdots,I_{N}\}$ and the corresponding short texts $\{T_{1},T_{2},\cdots,T_{N}\}$ are randomly sampled from the pre-training dataset. Each image and its corresponding text are treated as a positive pair, while others in the same batch are negative pairs. Then, the image encoder extracts image global feature $\mathbf{v}^{i}$ from the $i$-th image $I_{i}$ within the batch, while the text encoder obtains text feature $\mathbf{t}^{j}$ from the $j$-th text $T_{j}$. These two encoders are then optimized using contrastive loss $\mathcal{L}$, which consists of image-to-text loss $\mathcal{L}^{i2t}$ and text-to-image loss $\mathcal{L}^{t2i}$. It can be formulated as follows:
@@ 5.2 Ablation Studies
Table 2: Analyze the effectiveness of LoTLIP in language-image pre-training with long texts. The architecture of the image encoder is ViT-B/16. I2T and T2I indicate R@1 on text and image retrieval, respectively. We use 3M scale dataset for pre-training. “✓" indicates we add long texts in the training stage.
@@
I2T T2I I2T T2I I2T T2I I2TT2I Acc@1
LiT - 27.14 24.13 65.20 58.50 32.73 27.01 34.2024.07 43.76
LiT ✓ 47.96 44.92 84.97 81.70 73.66 66.73 43.5230.06 48.87
LiT+Long-CLIP* ✓ 47.21 46.61 83.82 83.01 74.60 67.49 43.78 26.74 46.82
LoTLIP49.4647.8284.9783.3376.4969.72 46.5631.5950.34
@@
I2T T2I I2T T2I I2T T2I I2T T2I Acc
0 - 47.96 44.92 84.97 81.7073.66 66.73 43.52 30.06 48.87
1 ✓ 49.57 46.55 84.97 82.6874.91 68.41 45.68 31.51 49.62
2 ✓ 49.46 47.82 84.97 83.33 76.49 69.7246.5631.59 50.34
3 ✓ 49.58 47.70 87.0984.3176.5170.20 46.48 31.60 50.36
4 ✓ 49.5848.30 86.76 84.1576.25 70.14 47.7031.8850.59
2 - 48.61 47.17 86.11 81.8676.14 69.31 47.70 31.34 49.88
2 ✓ 49.46 47.82 84.97 83.33 76.49 69.7246.56 31.59 50.34
Table 4: Zero-shot evaluation of different models on long-text-image retrieval tasks. I2T and T2I indicate R@1 on text and image retrieval, respectively.